\begin{thebibliography}{99}

\bibitem{vonNeumann55} J.\,von\,Neumann, {\it Mathematical Foundations of Quantum Mechanics} 
(Princeton University Press, Princeton, 1955).  

\bibitem{BZW71} {\it ``Quantum Theory and Beyond''}, ed. by T.\,Bastin (Cambridge U. Press, 
Cambridge, 1971).
%;     
%{\it ``Quantum Theory and Measurement''}, ed. by J.A.\,Wheeler and W.H.\,Zurek (Princeton U.  
%Press, Princeton, 1980).

\bibitem{Adler03} S.L.\,Adler, {\it Stud.\,Hist.\,Philos.\,Mod.\,Phys.} {\bf 34}, 135 (2003). 

\bibitem{Pearle} P.\,Pearle, {\it How Stands Collapse II}, arXiv: quant-ph/0611212 . 

\bibitem{Bassi} A.\,Bassi, {\it Dynamical Reduction Models: present status and future 
developments}, {\it J.\,Phys.: Conf. Series} (2007), to appear; arXiv: quant-ph/0701014 . 

\bibitem{I06} H.-T.\,Elze, {\it The Gauge Symmetry of the Third Kind and Quantum Mechanics as an Infrared Phenomenon}, Int.\,J.\,Qu.\,Inform. (2007), to appear; in:  
{\it A Sense of Beauty in Physics -- Festschrift for 
A.\,Di\,Giacomo}, ed. by M.\,D'Elia, K.\,Konishi, E.\,Meggiolaro and P.\,Rossi 
(Pisa Univ. Press, 2006), p.\,275, arXiv: quant-ph/0604142 .

\bibitem{Kibble78} T.W.B.\,Kibble, {\em Commun.\,Math.\,Phys.} {\bf 64}, 73 (1978). 

\bibitem{Kibble80} T.W.B.\,Kibble and S.\,Randjbar-Daemi, {\it J.\,Phys.\,A: Math.\,Gen.}  
{\bf 13}, 141 (1980).

\bibitem{Kiefer04} C.\,Kiefer, {\it Quantum Gravity} (Clarendon Press, Oxford, 2004). 

\bibitem{Iwo76} I.\,Bialynicki-Birula and J. Mycielski, {\em Ann.\,Phys.\,(N.Y.)} {\bf 100}, 
62 (1976). 

\bibitem{Weinberg89} S.\,Weinberg, {\em Phys.\,Rev.\,Lett.} {\bf 62}, 485 (1989); 
{\em Ann.\,Phys.\,(N.Y.)} {\bf 194}, 336 (1989).

\bibitem{DoebnerGoldin} H.-D.\,Doebner and G.A.\,Goldin, {\it Phys.\,Lett.} {\bf A162}, 397 
(1992); {\it Phys.\,Rev.} {\bf A54}, 3764 (1996); {\it Inst.\,Phys.\,Conf.\,Ser.} No\,{\bf 185}, 243 (2005). 

\bibitem{DoebnerGoldinNattermann} H.-D. Doebner, G.A. Goldin and P. Nattermann, 
{\it J.\,Math.\,Phys.} {\bf 40}, 49 (1999).  

\bibitem{Jackiw} R.\,Floreanini and R.\,Jackiw, {\it Phys.\,Rev.} {\bf D37}, 2206 (1988). 

%\bibitem{BarnesGhandour} T.\,Barnes and G.\,Ghandour, Nucl.\,Phys. {\bf B146}, 483 (1978). 
%\bibitem{Kiefer} C.\,Kiefer and A.\,Wipf, Ann.\,Phys. (NY) {\bf 236}, 241 (1994); 
%A.\,Duncan, H.\,Meyer-Ortmanns and R.\,Roskies, Phys.\,Rev. {\bf D36}, 3788 (1987).   
%C.\,Kiefer, Phys.\,Rev. {\bf D45}, 2044 (1992).  

\bibitem{Kiefer93} C.\,Kiefer, {\it Phys.\,Rev.} {\bf D47}, 5414 (1993). 
  
\bibitem{Barbour93} J.\,Barbour, {\it Phys.\,Rev.} {\bf D47}, 5422 (1993). 

\bibitem{tHooft01} G.\,'t\,Hooft, {\it J.\,Stat.\,Phys.} {\bf 53}, 323 (1988); 
{\it Class.\,Quant.\,Grav.} {\bf 16}, 3263 (1999); 
{\it Quantum Mechanics and Determinism}, in: Proc. of the Eighth Int. Conf. 
on ``Particles, Strings and Cosmology'', ed. by P.\,Frampton and J.\,Ng (Rinton Press, Princeton, 2001), p.\,275; 
{\it Determinism Beneath Quantum Mechanics}, arXiv: quant-ph/0212095 .

\bibitem{tHooft03} G.\,'t\,Hooft, {\it Int.\,J.\,Theor.\,Phys.} {\bf 42}, 355 (2003). 

\bibitem{tHooft06} G.\,'t\,Hooft, {\it The mathematical basis for deterministic quantum mechanics}, \\ arXiv: quant-ph/0604008 .

\bibitem{I05susy} H.-T.\,Elze, Phys.\,Lett. {\bf A335}, 258 (2005); {\it Braz.\,J.\,Phys.} 
{\bf 35}, 343 (2005); 
{\it A quantum field theory as emergent description of constrained supersymmetric classical dynamics}, 
Proceedings 8th Int. Conf. {\it ``Path Integrals. From Quantum Information to Cosmology''}, 
Prague, June 6-10, 2005, arXiv: hep-th/0508095 ; 
{\it J.\,Phys.: Conf.\,Ser.} {\bf 33}, 399 (2006). 

\bibitem{Vitiello01} M.\,Blasone, P.\,Jizba and G.\,Vitiello, {\it Phys.\,Lett.} {\bf A287}, 205 (2001);   
M.\,Blasone, E.\,Celeghini, P.\,Jizba and G.\,Vitiello, {\it Phys.\,Lett.} {\bf A310}, 393 (2003).  

\bibitem{ES02} H.-T.\,Elze and O.\,Schipper, {\it Phys.\,Rev.} {\bf D66}, 044020 (2002);   
H.-T.\,Elze, {\it Phys.\,Lett.} {\bf A310}, 110 (2003).  

%\bibitem{I04} H.-T. Elze, {\it Physica} {\bf A344}, 478 (2004); {\it Quantum Mechanics and 
%Discrete Time from ``Timeless'' Classical Dynamics}, in: \cite{E04}, p.\,196;  
%arXiv: quant-ph/0306096 . 
%\bibitem{E04} {\it ``Decoherence and Entropy in Complex Systems''}, ed. by H.-T.\,Elze, 
%Lecture Notes in Physics, Vol.\,633 (Springer-Verlag, Berlin Heidelberg New York, 2004). 

\bibitem{Blasone04} M.\,Blasone, P.\,Jizba and H.\,Kleinert, {\it Braz.\,J.\,Phys.} {\bf 35}, 497 (2005); {\it Phys.\,Rev.} {\bf A71}, 052507 (2005). 

\bibitem{Smolin} 
%L.\,Smolin, {\it Matrix Models as Non-Local Hidden 
%Variables Theories}, arXiv: hep-th/0201031 ;  
F.\,Markopoulou and L.\,Smolin, {\it Phys.\,Rev.} {\bf D70}, 124029 (2004). 

\bibitem{Adler} S.L.\,Adler, {\it ``Quantum Mechanics as an Emergent Phenomenon''}
%: The Statistical Dynamics of Global Unitary Invariant Matrix Models as the Precursors of Quantum Field Theory''} 
(Cambridge U. Press, Cambridge, 2005).

\bibitem{I05lambda} H.-T.\,Elze, {\it Quantum fields, cosmological constant and symmetry doubling}, {\it Int.\,J.\,Theor.\,Phys.} (2007), to appear, arXiv: hep-th/0510267 .

\bibitem{JackiwKerman} R.\,Jackiw and A. Kerman, {\em Phys.\,Lett.} {\bf A71}, 158 (1979); \\ 
H.-T. Elze, {\em Nucl.\,Phys.} {\bf B436}, 213 (1995); 
              {\em Phys.\,Lett.} {\bf B369}, 295 (1996); \\ arXiv: quant-ph/9710063 .

\bibitem{Ryder} L.H.\,Ryder, {\it ``Quantum Field Theory''} (Cambridge University Press, 
Cambridge, UK, 1986); \\ 
C.\,Itzykson and J.-B.\,Zuber, {\it ``Quantum Field Theory''} (McGraw-Hill, New York, 1985).  
  
\bibitem{Diosi84} L.\,Di\'osi, {\it Phys.\,Lett.} {\bf A105}, 199 (1984).  

\bibitem{Penrose98} R.\,Penrose, {\it Phil.\,Trans.\,R.\,Soc.} {\bf 356}, 1927 (1998). 

\bibitem{Diosi05} L.\,Di\'osi, {\it Braz.\,J.\,Phys.} {\bf 35}, 260 (2005). 

\bibitem{Carlip06} P.J.\,Salzman and S.\,Carlip, {\it A possible experimental test 
of quantized gravity}, \\ arXiv: gr-qc/0606120 . 

\bibitem{Diosi06} L.\,Di\'osi, {\it Notes on certain Newton gravity mechanisms of 
wave function localization and decoherence}, arXiv: quant-ph/0607110 . 

\end{thebibliography}

\documentclass[11pt,twoside]{article}
\begin{document}\hbadness=10000\thispagestyle{empty}
\title{A relativistic gauge theory of nonlinear quantum mechanics \\ 
and Newtonian gravity}

\author{Hans-Thomas Elze \\ \ \\ 
        Dipartimento di Fisica, Universit\`a di Pisa \\
        Largo Pontecorvo 3, I-56127 Pisa, Italia          \\ \ \\ 
        {\it E-mail: elze@df.unipi.it}}


\maketitle

\begin{abstract} A new kind of gauge theory is introduced, where the minimal 
coupling and corresponding covariant derivatives are defined in the space of functions 
pertaining to the functional Schr\"odinger picture of a given field theory. 
While, for simplicity, we study the example of a ${\cal U}(1)$ symmetry, 
this kind of gauge theory can accommodate other symmetries as well. 
We consider the resulting relativistic nonlinear extension of quantum mechanics and 
show that it incorporates gravity in the (0+1)-dimensional limit, similar to recently 
studied Schr\"odinger-Newton equations. Gravity is encoded here into a universal nonlinear extension of quantum theory. A probabilistic interpretation (Born's rule) holds, 
provided the underlying model is scale free. \\ 
{\it Keywords:} nonlinear functional Schr\"odinger equation, gauge symmetry, 
Newtonian gravity. \\    
{\it PACS:} 03.65.Ta, 03.70+k, 04.60.-m
\end{abstract}

\section{Introduction}
Linearity of the (functional) Schr\"odinger equation has been an essential and a 
puzzling ingredient of quantum (field) theory since its earliest days. Practically all  
physical phenomena show nonlinear behaviour when examined 
over a sufficiently large range of the dynamical parameters that determine an evolving 
object. What singles out the linear dynamics and validity of the superposition 
principle for the wave function(al)?
Quantum mechanics is very successfully tested experimentally under a 
wide range of laboratory conditions. Yet the mathematical structure of the 
theory, so far, hinges heavily on the linear structure embodied in linear 
operators acting on states represented by rays in a Hilbert space 
\cite{vonNeumann55,BZW71}.   

This raises the question: Are nonlinear extensions 
possible which agree with the standard formulation in its experimentally ascertained 
domain of validity? 

If so, could this alleviate the unresolved measurement problem 
\cite{vonNeumann55,BZW71,Adler03}? While the outcome of this second question 
is still open, 
it seems worth while to mention that in recent studies of the necessarily related 
wave function collapse or reduction mechanisms by Pearle \cite{Pearle} and 
by Bassi \cite{Bassi} the authors indicate that a nonlinear 
extension of quantum theory might ultimately account consistently for these effects. 

Our present     
aim is to report on a universal nonlinear extension of quantum field theory 
based on a new {\it functional gauge symmetry}, which operates 
on the space of field configurations rather than on the underlying spacetime \cite{I06}. 
In particular, we will argue that this theory essentially 
incorporates Newtonian gravity, which invites deliberation 
whether such an approach could be of wider use. 
Gravity, in this picture, appears as 
a manifestation of the nonlinearity of quantum mechanics. 
  
Among the numerous earlier works that have attempted to extend quantum theory in a nonlinear 
way, we should like to mention these: The work by Kibble and by Kibble and Randjbar-Daemi is 
close to ours in that they consider how nonlinear modifications of quantum field theory 
can be made compatible with Lorentz or more generally coordinate invariance 
\cite{Kibble78,Kibble80}. Besides considering a coupling of 
quantum fields to classical gravity according to general relativity, which induces an intrinsic nonlinearity \cite{Kibble80,Kiefer04}, these authors study mean-field type 
nonlinearities, where parameters of the model are state dependent through their 
assumed dependence on expectations of certain operators. The work by 
Bialynicki-Birula and Mycielski introduces a logarithmic nonlinearity into the nonrelativistic 
Schr\"odinger equation, which has the advantage that many of the nice features of 
quantum mechanics are left intact \cite{Iwo76}. A number of different nonrelativistic models 
of this kind have been systematically studied by Weinberg, offering also an assessment of 
the observational limits on such modifications of the Schr\"odinger equation 
\cite{Weinberg89}. 
  
Independently, Doebner and Goldin and collaborators have also studied nonlinear modifications 
of the nonrelativistic Schr\"odinger equation \cite{DoebnerGoldin,DoebnerGoldinNattermann}. 
While this was originally 
motivated by attempts to incorporate dissipative effects, they later have shown that classes  
of nonlinear Schr\"odinger equations, including many of those considered in the literature, 
for example, the one proposed in Ref.\,\cite{Iwo76}, 
can be obtained through nonlinear (in the wave function) transformations of the linear quantum 
mechanical equation. They coined the name ``gauge transformations of the third kind'' in this 
context, in analogy with the reasoning for gauge transformations of the second kind 
(corresponding to the usual minimal coupling). -- In distinction to their work, our functional 
gauge transformations, being set up for quantum field theory, work on the field configuration 
space over which the {\it wave functional} is defined. This will be most clearly recognized 
in the way we introduce {\it covariant} functional derivatives  
(cf. Eqs.\,(\ref{dtcov})--(\ref{dxcov}) in 
Section\,3). (Of course, the fact that functional derivatives come into play 
here is not new {\it per se}: they are to the functional Schr\"odinger picture of 
quantum field theory developed earlier by Jackiw and collaborators -- reviewed and 
generalized for fermions in \cite{Jackiw} --   
what ordinary derivatives are to quantum mechanics.)    

The necessity of generalizing quantum dynamics for quantum gravity has been discussed 
in view of the ``problem of time'' and the Wheeler-DeWitt equation 
by Kiefer and by Barbour \cite{Kiefer93,Barbour93}. -- We recall that this equation, 
playing the role of the Schr\"odinger equation there, is of 
the form of a constraint operator, i.e. the Hamiltonian of canonical gravity, 
acting on the wave functional, $\hat H\Psi =0$. There are two unpleasant features: 
no time derivative appears \cite{Kiefer04,Kiefer93} and, since $\hat H$ is hermitean, there seems to be no indication of complex solutions \cite{Barbour93}! -- 
Therefore, both authors pointed out that nonlinear 
modifications would be a wellcome remedy and in Ref.\,\cite{Kiefer93} it was proposed that 
these may assume the form of a ``supergauge potential'' defined on configuration space. 
While formally analogous to the gauge connection in the covariant derivatives introduced  
here, only some preliminary interpretation has been offered that such connection 
might effectively represent certain quantum (vacuum) effects of matter. 

In distinction, based on the proposed functional gauge symmetry, all dynamical and 
constraint equations of our theory are consistently derived from  
a gauge and Lorentz invariant action, as we shall discuss. From the outset, this  
has nothing to do with gravity, in particular, but may be applied to any  
quantum field theory. 

The importance of maintaining a probabilistic 
interpretation of the wave function following the Born rule is stressed in all 
previous works. We will recover this as well.  
However, no understanding of the origins of the proposed nonlinearities has been provided, 
except in the obvious case of gravity coupling studied by Kibble and Randjbar-Daemi 
\cite{Kibble80}.  
Presently, this is achieved by the gauge principle. Furthermore, 
we obtain the surprising result that our theory automatically 
incorporates gravity in its simplest Newtonian form, which 
will be discussed in Section\,6. 

Part of the motivation for the present work comes from recent considerations of the 
possibility of a deterministic foundation of quantum mechanics, as it has already been 
verified in a number of models 
\cite{tHooft01,tHooft03,tHooft06,I05susy,Vitiello01,ES02,Blasone04,Smolin,Adler,I05lambda}. 
While general principles and physical mechanisms ruling the construction of a deterministic 
classical model underlying a given quantum field theory are hard to come by, cf. 
Ref.\,\cite{tHooft06}, the already known models are quite promising, amounting to an existence 
proof that the quantum harmonic oscillator can be understood completely in classical 
deterministic terms, see Refs.\,\cite{tHooft01,tHooft03,ES02}. 

In general, we expect that 
with improved understanding of the emergence of quantum mechanics also resulting nonlinear 
corrections to quantum mechanics, as it is, should become visible. Note that any model 
which has an evolution equation that is linear in the wave function, i.e. without any 
nonlinear feedback, can always be cast into the form of a Schr\"odinger equation, possibly 
with modified potentials etc. Nonlinearity seems unavoidable, if one wants to go beyond 
the canonical framework of quantum theory. In the present work, such a nonlinear 
extension of quantum field theory is a central aspect.  

The paper is organized as follows. 
In Section\,2, we recapitulate the work of Ref.\,\cite{Kibble80} which forms the basis 
for our argument that the gauge invariant (quantum) action introduced in Section\,3 
can be written in a Lorentz invariant way, despite the presence of a fundamental length 
parameter. In Section\,4, the dynamical and constraint equations are presented and 
a crucial ``nonlinearity factor'' of the action is determined. Section\,5 is dedicated 
to the discussion of the validity of the Born rule in the resulting nonlinear quantum theory.  
In Section\,6, it is demonstrated that it leads to the Schr\"odinger-Newton equations in 
the onedimensional limit considering stationary states. 
Concluding remarks follow in Section\,7.      

\section{The Schr\"odinger picture for given background space-time}
Following the work of Kibble and Randjbar-Daemi \cite{Kibble80}, 
we consider a four-dimensional globally hyperbolic manifold ${\cal M}$ with a 
{\it given} metric $g_{\mu\nu}$ of signature $(1,-1,-1,-1)$.\footnote{Units 
are chosen such that $\hbar =c=1$.} Then, it is 
always possible to introduce a global slicing into space-like hypersurfaces, 
such that a chosen family of such surfaces, $\{\sigma (t)\}$, is locally determined 
by: 
\begin{equation}\label{xmu}
x^\mu =x^\mu (\xi^1,\xi^2,\xi^3;t) 
\;\;, \end{equation} 
in terms of intrinsic coordinates $\xi^r$, and there exists an everywhere 
time-like vectorfield $n^\mu$, the normal, with $n_\mu n^\mu =1$ and 
$n_\mu x^\mu_{,r}=0$, where $x^\mu_{,r}\equiv\partial x^\mu/\partial\xi^r$. 
We will make use of the derivative with respect to $t$ at fixed $\xi^r$ of a  
function $f$, $\dot f\equiv\partial f/\partial t\vert_{\xi}$. In particular, 
then, the lapse function $N$ and shift vector $N^r$ are introduced through 
the relation $\dot x^\mu =Nn^\mu +N^rx^\mu_{,r}$, the geometrical meaning 
of which is illustrated, for example, in Chapter\,3.3 of reference \cite{Kiefer04}. 

We assume a given Lagrangean $L$ of a field theory, 
such as for a real scalar field $\phi$: 
\begin{equation}\label{Lagrangean} 
L\equiv\frac{1}{2}g^{\mu\nu}\partial_\mu\phi\partial_\nu\phi -V(\phi )
\;\;, \end{equation} 
where $V(\phi )$ incorporates mass or selfinteraction terms. Then, 
the invariant action is defined by: 
\begin{equation}\label{action} 
S\equiv\int\mbox{d}^4x\;\sqrt{-g}L 
\;\;, \end{equation} 
where $g\equiv\det g_{\mu\nu}$. This, in turn, yields the stress-energy tensor $T^{\mu\nu}$  
through the relation: 
\begin{equation}\label{Tmunu} 
\frac{1}{2}\sqrt{-g}T^{\mu\nu}\equiv\frac{\delta S}{\delta g_{\mu\nu}} 
=\frac{1}{2}\sqrt{-g}\left (\partial^\mu\phi\partial^\nu\phi -g^{\mu\nu}L\right )
\;\;. \end{equation} 
With the help of the induced metric $\gamma_{rs}$ on $\sigma (t)$, 
$\gamma_{rs}\equiv g_{\mu\nu}x^\mu_{,r}x^\nu_{,s}$, and the hypersurface 
element $\mbox{d}\sigma_\mu\equiv\mbox{d}^3\xi\;\sqrt{-\gamma}n_\mu$,  
the surface-dependent Hamiltonian can be defined: 
\begin{equation}\label{Hamiltonian} 
H(t)\equiv\int_{\sigma (t)}\mbox{d}\sigma_\mu T^\mu_\nu\dot x^\nu 
\;\;. \end{equation} 
In the simplest case, with $\dot x^\mu =\delta^\mu_0$ (i.e., $N=1$, $N^r=0$) and 
$x^\mu_{,r}=\delta^\mu_r$, these relations reduce to $\gamma_{rs}=g_{rs}$ and 
$H(t)=\int_{\sigma (t)}\mbox{d}^3\xi\;T^{00}$, as expected.  
  
If the stress-energy tensor can be expressed in terms of 
canonical coordinates and momenta, for example, the scalar field $\phi =\phi (\xi^i,t)$ and 
its conjugated momentum $\Pi =\Pi (\xi^i,t)$ on time slices $\sigma (t)$, we assume that the corresponding quantized theory exists, with $\phi$ and $\Pi$ fullfilling the usual   
equal-$t$ commutation relation. 
Of course, matters are not that simple in a general curved background.  
Therefore, a heuristic 
derivation of the Schr\"odinger picture from the manifestly covariant Heisenberg 
picture has been presented in Ref.\,\cite{Kibble80}. We will not pursue this further, since  
our aim here is simply to recover their Lorentz invariant form of 
the functional Schr\"odinger equation, a generalization of which will follow 
from the action principle to be considered in the course of this 
work.\footnote{As remarked in Ref.\,\cite{Kibble80}, 
the derivation from an action principle will guarantee the general coordinate 
invariance of the theory. However, the 
Schr\"odinger picture clearly depends on the slicing of space-time as well 
as on the parametrisation of the slices. Thus,   
invariance under surface deformations 
-- which can be restricted to diffeomorphism invariance \cite{Kiefer04} -- 
is not implied here.}    

In any case, the functional Sch\"odinger 
equation obtained by Kibble and Randjbar-Daemi appears naturally 
as one would guess ($\hbar =c=1$): 
\begin{equation}\label{Schroedinger} 
i\dot\Psi =H(t)\Psi 
\;\;. \end{equation} 
Using the surface element $\mbox{d}\sigma_\mu$ given above, together 
with Eq.\,(\ref{Hamiltonian}), and: 
\begin{equation}\label{tDerivLocal}
\dot\Psi = \int_{\sigma (t)}\mbox{d}^3\xi\;\dot x^\mu\frac{\delta}{\delta x^\mu}\Psi
\;\;, \end{equation} 
the Schr\"odinger equation can also be 
represented in a local form: 
\begin{equation}\label{SchroedingerLocal} 
i\frac{\delta}{\delta x^\mu}\Psi =\sqrt{-\gamma}n_\nu T^\nu_\mu\Psi
\;\;. \end{equation}  
We take from this section that the functional Schr\"odinger equation 
can be written in a way that makes the behaviour under Lorentz transformations  
explicit. This applies, in particular, to the case of a flat background space-time, 
where field quantization is well understood.    

\section{The gauge invariant action}
We consider the generic scalar field theory described by the Lagrangean of 
Eq.\,(\ref{Lagrangean}), 
while internal symmetries and fermions can be introduced as we discussed earlier 
in the second of Refs.\,\cite{I06}.  
Furthermore, specializing the result of the previous section for {\it Minkowski space}, 
we find: 
\begin{equation}\label{HamiltonianMinkowski} 
H(t)=\int_{\sigma (t)}\mbox{d}^3\xi\;T^{00}
=\int\mbox{d}^3x\;\Big\{ -\frac{1}{2}\frac{\delta^2}{\delta\phi^2}
+\frac{1}{2}(\nabla\phi )^2+V(\phi )\Big\}
\equiv H[\hat\pi ,\phi]
\;\;, \end{equation} 
i.e., the usual Hamiltonian which is independent of the parameter time $t$;   
intrinsic and Minkowski space coordinates have been identified, $\vec\xi =\vec x$. 
   
Here, in the Hamiltonian of Eq.\,(\ref{HamiltonianMinkowski}) in particular, 
quantization is implemented by substituting the canonical momentum $\pi$ 
conjugate to the field $\phi$ (playing the role of ``coordinate''): 
\begin{equation}\label{momentum}
\pi (\vec x)\;\longrightarrow\;\hat\pi (\vec x)\equiv
\frac{1}{i}\frac{\delta}{\delta\phi (\vec x)}
\;\;. \end{equation}
Correspondingly, we have $\Psi =\Psi [\phi ;t]$, i.e. 
a time dependent functional, in this coordinate representation, and $\dot\Psi =\partial_t\Psi$. 
So far, this is the usual functional Schr\"odinger picture of quantum field theory 
applied to the chosen example of a scalar model \cite{Jackiw,JackiwKerman,Ryder}.  

Next, we introduce {\it functional gauge transformations} \cite{I06}:  
\begin{equation}\label{localfuncgt} 
\Psi'[\phi ;t]=\exp (i\Lambda [\phi ;t])\Psi [\phi ;t]
\;\;, \end{equation} 
where $\Lambda$ denotes a time dependent real functional. These ${\cal U}$(1) 
transformations are local in the space of field configurations.   
They differ from the usual    
gauge transformations in QFT, since we 
introduce covariant derivatives by the following replacements:  
\begin{eqnarray}\label{dtcov}
\partial_t&\longrightarrow &{\cal D}_t\equiv
\partial_t-i{\cal A}_t[\phi ;t]
\;\;, \\ \label{dxcov} 
\frac{\delta}{\delta\phi (\vec x)}&\longrightarrow &
{\cal D}_{\phi (\vec x)}\equiv
\frac{\delta}{\delta\phi (\vec x)}-i{\cal A}_\phi [\phi ;t,\vec x]
\;\;. \end{eqnarray} 
The real functional ${\cal A}$ presents a new kind of `potential' or `connection'. 
Generally, ${\cal A}$ depends on $t$. However, 
it is a {\it functional} of $\phi$ in Eq.\,(\ref{dtcov}), while it is a 
{\it functional field} in Eq.\,(\ref{dxcov}). 
We distinguish these components of ${\cal A}$ by the subscripts. 
Furthermore, the `potentials' are required to transform as: 
\begin{eqnarray}\label{Afunctionalgt} 
{\cal A}'_t[\phi ;t]&=&{\cal A}_t[\phi ;t]+
\partial_t\Lambda [\phi ;t]
\;\;, \\ \label{Afunctiongt}  
{\cal A}'_\phi [\phi ;t,\vec x]&=&
{\cal A}_\phi [\phi ;t,\vec x]+
\frac{\delta}{\delta\phi (\vec x)}\Lambda [\phi ;t]
\;\;. \end{eqnarray} 
Applying Eqs.\,(\ref{localfuncgt})--(\ref{Afunctiongt}), it follows 
that the correspondingly generalized functional Schr\"odinger equation 
is invariant under the ${\cal U}$(1) gauge transformations. 

Furthermore, it is suggestive to introduce an invariant `field strength':
\begin{equation}\label{field} 
{\cal F}_{t\phi}[\phi ;t,\vec x]\equiv 
\partial_t{\cal A}_\phi [\phi ;t,\vec x]
-\frac{\delta}{\delta\phi (\vec x)}
{\cal A}_t[\phi ;t]
\;\;, \end{equation} 
in close analogy to ordinary gauge theories; 
note that ${\cal F}_{t\phi}=[{\cal D}_t,{\cal D}_\phi]/(-i)$. 
     
We now postulate a consistent dynamics for the gauge `potential' 
${\cal A}$, in order to give a meaning to the above `minimal coupling' 
prescription. All elementary fields supposedly are   
present as the coordinates on which the wave 
functional depends -- presently just a scalar field, besides time.  
We consider the following ${\cal U}$(1) invariant action:  
\begin{equation}\label{Action}
\Gamma\equiv\int\mbox{d}t\mbox{D}\phi\;\Big\{ 
\Psi^*\Big ({\cal N}(\rho )
\stackrel{\leftrightarrow}{i{\cal D}}_t
-H[\frac{1}{i}{\cal D}_{\phi},\phi ]\Big )\Psi 
+\frac{l^2}{2}\int\mbox{d}^3x\;\big ({\cal F}_{t\phi}
\big )^2\Big\}
\;\;, \end{equation} 
where    
$\Psi^*{\cal N}\stackrel{\leftrightarrow}{i{\cal D}}_t\Psi
\equiv\frac{1}{2}{\cal N}
\{\Psi^*i{\cal D}_t\Psi
+(i{\cal D}_t\Psi )^*\Psi\}$, and with   
a dimensionless real function ${\cal N}$ which depends on the density:  
\begin{equation}\label{rho}
\rho [\phi ;t]\equiv\Psi^*[\phi ;t]\Psi [\phi ;t]
\;\;. \end{equation}
We shall see shortly that ${\cal N}$ incorporates a \underline{necessary} {\it nonlinearity}; it 
will be uniquely determined in Section\,4, cf. Eq.\,(\ref{nonlin}). The fundamental parameter $l$ has dimension $[l]=[length]$, for dimensionless measure $\mbox{D}\phi$ and $\Psi$,  
independently of the dimension of space-time.\footnote{Note that this parameter has \underline{necessarily} 
the dimension of a length, in order to give the action its correct dimension; it presents the 
coupling constant of our theory and will be related to Newton's constant in Section\,6. By suitably  
rescaling the gauge `potentials', the coupling constant could be moved to the covariant derivatives, as originally 
discussed \cite{I06}; however, as is familiar from ordinary non-Abelian gauge theories, 
the equivalent action is often more convenient where the coupling appears only in one place.} 

The action $\Gamma$ generalizes the 
action for the wave functional of a scalar field, which has been 
employed for applications of Dirac's 
variational principle to QFT, e.g., in Refs.\,\cite{Kibble80,JackiwKerman}. 
The quadratic part in ${\cal F}_{t\phi}$ is the 
simplest possible extension, i.e. local in  
$\phi$ and quadratic in the derivatives, together with 
the nonlinearity ${\cal N}(\rho )$ introduced here. 

An immediate consequence of the ${\cal U}$(1) invariance is 
that the Hamiltonian $H$, unlike in QFT, 
cannot be arbitrarily shifted by a constant $\Delta E$, gauge transforming  
$\Psi\rightarrow\exp (-i\Delta Et)\Psi$. Thus, there is an absolute 
meaning to the zero of energy in this theory. 

Translation invariance of the action, Eq.\,(\ref{Action}), 
gives rise to a {\it conserved energy functional}, where a contribution 
which is solely due to 
${\cal A}_t$ and ${\cal A}_\phi$ is added to the matter term, 
which is modified by the covariant derivatives. 

Furthermore, following from the discussion of Section\,2, 
the Lorentz invariance 
of this theory is guaranteed. In particular, the action can be written 
in a Lorentz (and Poincar\'e) invariant way, using the appropriate surface-dependent 
Hamiltonian, cf. Eq.\,(\ref{Hamiltonian}), despite that a fundamental length $l$ enters 
here.\footnote{The coordinates $x^\mu$, of course, must not be confused with 
the intrinsic coordinates $\xi^i$ and time parameter $t$.} 

The action depends on 
$\Psi ,\Psi^*,{\cal A}_t$, and 
${\cal A}_\phi$ separately. 
While a Hamiltonian formulation is possible, the equations of motion and a constraint 
can be obtained directly by 
varying $\Gamma$ with respect to these variables. 

\section{The dynamical and constraint equations}
The dynamical equations of motion were previously obtained in 
Refs.\,\cite{I06} and are reproduced here 
for convenience. The gauge covariant equation for the $\Psi$-functional is:   
\begin{equation}\label{ginvfSscalar}
\left (\rho {\cal N}(\rho )\right )' 
i{\cal D}_t\Psi [\phi ;t]
=H[\frac{1}{i}{\cal D}_{\phi},\phi ]
\Psi [\phi ;t]
\;\;, \end{equation}
where $f'(\rho )\equiv\mbox{d}f(\rho )/\mbox{d}\rho$. This replaces 
the usual functional Schr\"odinger equation (and similarly its adjoint).  

The nonlinear Eq.\,(\ref{ginvfSscalar}) preserves the normalization of 
$\Psi$. Fixing it at an initial parameter time, 
in terms of an arbitrary constant ${\cal C}_0$:  
\begin{equation}\label{norm}
\langle\Psi |\Psi\rangle\equiv\int\mbox{D}\phi\;\Psi^*\Psi = {\cal C}_0 
\;\;, \end{equation} 
it is conserved under further evolution, while  
the overlap of two different states, $\langle\Psi_1|\Psi_2\rangle$, may 
vary. This is indicative of the standard probability interpretation 
related to $\Psi^*\Psi$, which we will discuss in the next 
section in more detail. 

Next, there is an invariant `gauge field equation': 
\begin{equation}\label{fieldeq} 
\partial_t{\cal F}_{t\phi}[\phi;t,\vec x]
=-\frac{1}{2il^2}\left ( 
\Psi^*[\phi;t]{\cal D}_{\phi (\vec x)}\Psi [\phi;t]
-\Psi [\phi;t]({\cal D}_{\phi (\vec x)}\Psi [\phi;t])^*
\right )
\;\;, \end{equation}
which completes the dynamical equations.

However, there is no time derivative acting on the variable 
${\cal A}_t$ in the action. Therefore, it acts as a Lagrange multiplier for a constraint, 
which is the gauge invariant `Gauss' law':
\begin{equation}\label{Gauss}
\int\mbox{d}^3x\;\frac{\delta}{\delta\phi (\vec x)}{\cal F}_{t\phi}[\phi ;t,\vec x]
=-\frac{1}{l^2}\rho {\cal N}(\rho )
\;\;. \end{equation} 
Of course, it differs from the usual one in QED, for example. 
This raises the 
question, whether our functional ${\cal U}$(1) gauge symmetry is compatible 
with the presence of standard internal symmetries. This is answered  
affirmatively in the second of Refs.\,\cite{I06}. 

The Eq.\,(\ref{Gauss}) can be combined with Eq.\,(\ref{fieldeq}) to result in a 
continuity equation: 
\begin{equation}\label{continuity}
0=\partial_t\Big (\rho{\cal N}(\rho )\Big )
-\frac{1}{2i}\int\mbox{d}^3x\;\frac{\delta}{\delta\phi (\vec x)}
\left ( 
\Psi^*{\cal D}_{\phi (\vec x)}\Psi 
-\Psi ({\cal D}_{\phi (\vec x)}\Psi )^*
\right )
\;, \end{equation}
expressing local ${\cal U}$(1) `charge' conservation in the space of field configurations. 
Functionally integrating Eq.\,(\ref{Gauss}), we find that the 
total `charge' $Q$ has to vanish at all times: 
\begin{equation}\label{charge}
Q(t)\equiv\frac{1}{l^2}\int\mbox{D}\phi\;\rho{\cal N}(\rho )=0
\;\;, \end{equation}    
since the functional integral of a total derivative is zero \cite{Ryder}. This 
is different from integrating the usual Gauss' law in electrodynamics over all 
space, for example, were there can be a flux of the fields out to infinity. -- 
The necessity of the nonlinearity now becomes obvious. 
Without it, the vanishing total `charge' could not be implemented, as it 
would be in conflict with the normalization, Eq.\,(\ref{norm}).  

We proceed to determine the nonlinearity factor, 
${\cal N}(\rho )\neq 1$.
In fact, we would like to implement Eq.\,(\ref{charge}), similarly as the 
normalization, at an initial parameter time $t$. Since it has to be a 
constant of motion, $\partial_tQ(t)=0$, we express this, 
with the help of Eq.\,(\ref{ginvfSscalar}), 
as a condition on $\rho{\cal N}(\rho )$. It is easily seen that the 
only solution here is a linear function: 
\begin{equation}\label{nonlin}
\rho {\cal N}(\rho )={\cal C}_1\Big (\rho -{\cal C}_0(\int\mbox{D}\phi )^{-1}\Big ) 
\;\;, \end{equation}
if one wants to avoid further constraining $\Psi$ or $\Psi^{*}$; 
the latter would make it more difficult, if not impossible, to obtain linear quantum mechanics 
as a limiting case.\footnote{In Ref.\,\cite{I06}, a logarithmic form was chosen.    
In view of the present discussion, however, it should be dismissed, since it is not 
based on a constant of motion.}  

Evidently, the volume of the space of fields, $\Omega\equiv\int\mbox{D}\phi$, needs to be 
regularized, as well as the second functional derivatives at coinciding points 
which appear. A cut-off on field amplitudes has to be introduced together, for 
example, with dimensional regularization \cite{Ryder} or, more convenient 
here, the point-splitting technique \cite{Jackiw}. Clearly, a related renormalization 
procedure is an interesting subject for further study, since it has to take into 
account the new functional gauge symmetry.

\section{Probability versus `charge'}  
The {\it homogeneity property} is necessary for the probability interpretation 
of the density $\rho =\Psi^*\Psi$ according to Born's rule 
\cite{Kibble78,Iwo76,Weinberg89}: 
$\Psi$ and $z\Psi$ ($z\in\mathbf{Z}$) have to present the same   
physical state. In this way, states are associated 
with rays in a Hilbert space (instead of vectors). 

For the present case, it is useful to consider the 
set of scale transformations: 
\begin{equation}\label{scale1} 
\rho ={\cal C}_0^{\;a}{\cal C}_1^{-1}\rho'\;\;,\;\;\;
\int\mbox{D}\phi ={\cal C}_0^{1-a}{\cal C}_1\int\mbox{D}\phi'
\;\;, \end{equation} 
such that $\int\mbox{D}\phi'\rho'=1$; we recall that the real measure 
$\mbox{D}\phi$ and constants ${\cal C}_{0,1}$ are chosen dimensionless, without loss 
of generality; $a$ is real. Furthermore, we rescale: 
\begin{equation}\label{scale2}
(\vec x;t)={\cal C}_0^{-a/2}{\cal C}_1^{1/2}(\vec x';t')\;\;,\;\;\; 
(\phi ;{\cal A}_t)={\cal C}_0^{a/2}{\cal C}_1^{-1/2}(\phi';{\cal A}'_t)
\;\;, \end{equation} 
and, consistently: 
\begin{equation}\label{scale3} 
(\delta_\phi ;{\cal A}_\phi ) =
{\cal C}_0^a{\cal C}_1^{-1}(\delta_{\phi'};{\cal A}_{\phi'})  
\;\;. \end{equation}   
Under these transformations, the action transforms as: 
\begin{equation}\label{Actionscale} 
\Gamma ={\cal C}_1\Gamma' 
\;\;, \end{equation}  
where $\Gamma'$ is defined like $\Gamma$, Eq.\,(\ref{Action}), however, 
replacing all quantities by the primed ones.   
One arrives at this result, provided the Hamiltonian $H$, cf. Eq.\,(\ref{HamiltonianMinkowski}),  
contains {\it no dimensionfull constants}, such as in a Lagrangean mass term, $\propto m^2\phi^2$, 
which could be contained in our model; a selfinteraction of the 
form $\propto\lambda\phi^4$ introduces a dimensionless coupling $\lambda$ instead.   

There are several implications. -- First, the scale transformations change the overall scale 
of the action, 
say, in units of $\hbar$, by the constant factor ${\cal C}_1$. This is equivalent to the  
rescaling $\hbar =\hbar'/{\cal C}_1$. However, since we prefer to choose units such that 
$\hbar =1$, we should also fix ${\cal C}_1=1$, henceforth.\footnote{The overall sign of 
$\rho {\cal N}(\rho )$, cf. Eq.\,(\ref{nonlin}), is chosen with hindsight, see Section\,6.} -- Second, since the 
constant ${\cal C}_0$ does not affect the transformation of $\Gamma$, we can always 
choose to normalize the wave functional to ${\cal C}_0=1$, see Eq.\,(\ref{norm}).  

We see that states, as far as $\Psi$ is concerned, are represented by rays. Therefore, 
a probability interpretation of $\Psi^*\Psi$ according to the {\it Born rule} can be maintained.  
This is in agreement with the obervation that Eq.\,(\ref{ginvfSscalar}), if it were 
not for the presence of the covariant derivatives, now appears like the usual 
functional Schr\"odinger equation. Summarizing the previous discussion, we now have: 
\begin{eqnarray}\label{nonlin1}
\rho {\cal N}(\rho )&=&\rho -(\int\mbox{D}\phi )^{-1}
\;\;, \\ [1ex] \label{ginvfSscalar1}   
i{\cal D}_t\Psi [\phi ;t]
&=&H[\frac{1}{i}{\cal D}_{\phi},\phi ]
\Psi [\phi ;t]
\;\;. \end{eqnarray}
However, it must be stressed that the `potentials' ${\cal A}_t$ and ${\cal A}_\phi$ 
are selfconsistently determined through Eqs.\,(\ref{fieldeq})--(\ref{Gauss}). 
Therefore, we arrive here at intrinsically {\it nonlinear quantum 
mechanics}.\footnote{In the second of Refs.\,\cite{I06}, we have argued that 
{\it microcausality} of the present theory holds. -- The {\it weak superposition 
principle} \cite{Iwo76}, generally, must be expected to fail: 
for two non-overlapping sources adding to the right-hand sides of 
Eqs.\,(\ref{fieldeq})--(\ref{Gauss}), the resulting `potentials' must be expected 
to propagate away from the sources in field space. Thus, the sum of two non-overlapping 
solutions $\Psi_{1,2}$ will hardly present a solution of the coupled equations. 
However, {\it if} two {\it stationary non-overlapping solutions} exist, {\it then} their sum 
also presents a solution; see the stationary equations in Section\,6.}     

The difference to standard quantum mechanics also shows up clearly in Eq.\,(\ref{continuity}), 
with the first term now replaced by $\partial_t\rho$: the flux of probability over 
the space of field configurations is affected nonlinearly by $\Psi^*$ and $\Psi$ 
through the `potentials'.  

Finally, we remark that {\it in presence of dimensionfull parameters} in the 
Hamiltonian the above scale symmetry, Eqs.\,(\ref{scale1})--(\ref{Actionscale}), 
breaks down. In particular, then the normalization of $\Psi$ cannot be chosen 
freely; correspondingly, rays break into inequivalent vectors. In this situation, 
it is appropriate to consider $\Psi$ and 
$\Psi^*$ as giving rise to two oppositely `charged' real components of the wave functional, 
$\Psi_+\equiv (\Psi +\Psi^*)/\sqrt 2$ and $\Psi_-\equiv (\Psi -\Psi^*)/i\sqrt 2$, which 
interact, while preserving the normalization of $\Psi^*\Psi$. Different normalizations, 
then, correspond to physically different sectors of the theory, i.e. a charge superselection 
rule.    

However, the {\it absence of the homogeneity property} modifies the usual 
measurement theory. In particular, the usual ``reduction of the wave packet'' 
postulate \cite{vonNeumann55} cannot be maintained, in this case. This has been discussed in 
detail in Ref.\,\cite{Kibble78} and formed the starting point for the particular nonlinear theory proposed there, mentioned before in Section\,1. 

\section{Stationary states and the Schr\"odinger-Newton equations}
The time dependence in  
Eqs.\,(\ref{ginvfSscalar})--(\ref{Gauss}) can be separated with the Ansatz    
$\Psi [\phi ;t]\equiv\mbox{exp}(-i\omega t)\Psi_\omega [\phi ]$, 
$\omega\in\mathbf{R}$, and consistently 
assuming {\it time independent} ${\cal A}$-functionals. Thus, 
the Eq.\,(\ref{ginvfSscalar}), together with Eq.\,(\ref{nonlin1}), yields: 
\begin{equation}\label{PsiZeroEq} 
\omega\Psi_\omega [\phi ]
=H[\frac{1}{i}{\cal D}_{\phi},\phi ]\Psi_\omega [\phi ]-{\cal A}_t[\phi ]
\Psi_\omega [\phi ]
\;\;, \end{equation}
with ${\cal D}_\phi = \frac{\delta}{\delta\phi}+i{\cal A}_\phi$ and  
$\rho_\omega\equiv\Psi_\omega^*[\phi ]\Psi_\omega [\phi ]$. 
From Eq.\,(\ref{fieldeq}) follows: 
\begin{equation}\label{fieldeq0} 
\frac{1}{2i}\left ( 
\Psi^*_\omega [\phi ]{\cal D}_{\phi (\vec x)}\Psi_\omega [\phi ]
-\Psi_\omega [\phi ]({\cal D}_{\phi (\vec x)}\Psi_\omega [\phi ])^*
\right )=0
\;\;, \end{equation}
which expresses the vanishing of the `current' in the stationary 
situation. -- 
Applying a time independent gauge transformation, 
cf. Eqs.\,(\ref{localfuncgt}), (\ref{Afunctiongt}), the stationary wave functional 
can be made {\it real}. Then, the Eq.\,(\ref{fieldeq0}) implies ${\cal A}_\phi=0$;   
consequently, ${\cal D}_{\phi}\rightarrow\frac{\delta}{\delta\phi}$ 
everywhere. Finally, `Gauss' law', Eq.\,(\ref{Gauss}), 
determines ${\cal A}_t$: 
\begin{equation}\label{Gauss0}  
\int\mbox{d}^3x\;\frac{\delta^2}{\delta\phi (\vec x)^2}{\cal A}_t[\phi ]
=\frac{1}{l^2}\Big (\rho -(\int\mbox{D}\phi )^{-1}\Big )
\;\;, \end{equation} 
which has to be solved selfconsistently together with Eq.\,(\ref{PsiZeroEq}). -- 
Separation of the time dependence thus leads to two coupled 
equations. They represent a field theoretic generalization 
of the stationary Schr\"odinger-Newton equations, as we shall explain. 

The time dependent {\it Schr\"odinger-Newton equations} for a particle of mass $m$ are given by: 
\begin{equation}\label{SN} 
i\hbar\partial_t\psi =-\frac{\hbar^2}{2m}\nabla^2\psi -m\Phi\psi\;\;,\;\;\; 
\nabla^2\Phi =4\pi Gm\vert\psi\vert^2 
\;\;, \end{equation}
where $G\equiv l_P^2c^2/\hbar$ is Newton's gravitational constant 
(here related to the Planck length $l_P$) and $\Phi$ denotes the gravitational 
potential. They represent the nonrelativistic approximation to ``semiclassical gravity'', 
i.e. Einstein's field equations coupled to the expectation value of the operator-valued stress-energy tensor of quantum matter -- see, for example, the 
Refs.\,\cite{Kibble80,Diosi84,Penrose98,Diosi05,Carlip06,Diosi06}, and further references 
therein. They play an important role in arguments related to ``semiclassical gravity'', 
to gravitational self-localization of mesoscopic or macroscopic mass distributions, and 
to the role of gravity in Di\'osi's and Penrose's objective reduction scenarios. 
  
Considering a Universe which consists only of a single point, 
we find that our field theory equations (\ref{PsiZeroEq}) and (\ref{Gauss0}) 
reduce to the stationary 
Schr\"odinger-Newton equations {\it in one dimension}. Appropriate rescalings by powers 
of $l$, $m$, $\hbar$, and $c$ of the various quantities have to be incorporated, 
in order to give the 
equations their onedimensional form, where Newton's constant $G$ is a dimensionless 
parameter. If there is a nonzero potential $V(\phi )$ in our Hamiltonian, 
this extends the Schr\"odinger equation in (\ref{SN}) by an additional potential term. --  
Explicitly, keeping units such that $\hbar =c=1$ and considering the Hamiltonian of 
Eq.\,(\ref{HamiltonianMinkowski}) with $V(\phi )\equiv 0$, 
the following substitutions have to be performed, in order to arrive at 
the stationary limit of Eqs.\,(\ref{SN}): 
\begin{eqnarray}\label{subs1}
|\Psi |^2\;\longrightarrow\;4\pi G^2m|\psi |^2&,&\; {\cal A}_t\;\longrightarrow\; m\Phi\;\;, 
\\ [1ex] \label{subs2} 
\int\mbox{d}^3x\;\frac{\delta^2}{\delta \phi (x)^2}\;\longrightarrow\;\frac{1}{m}
\frac{\mbox{d}^2}{\mbox{d}q^2}&,&\;\int{\cal D}\phi\;\longrightarrow\;(4\pi G^2m)^{-1}
\int_{-Q/2}^{+Q/2}dq=Q/4\pi G^2m
\;\;, \end{eqnarray} 
where $m$ is the relevant (particle) mass scale and Q denotes a regulator length, much larger 
than any length scale of the one-dimensional system. Of course, the gradient terms of the 
Hamiltonian, $\propto (\nabla\phi )^2$, do not contribute in this limit (``a single point has 
no neighbours'').     


It seems remarkable that the gravitational interaction arises here in the space of 
quantum states (configuration space). Yet, in view of the fundamental length $l$ present 
in the action, Eq.\,(\ref{Action}), it is perhaps not a complete surprise that our gauge 
theory incorporates gravity. We notice, however, also a {\it deviation from 
Newtonian gravity}, presented by the constant term on the right-hand side of Eq.\,(\ref{Gauss0}). 
While it is natural to let this term become arbitrarily small in the quantum mechanical limit 
just discussed, its presence was shown to be necessary for the full theory in Section\,4.  
This is an important topic for further study, related to the regularization of the theory. 


In Ref.\,\cite{Carlip06}, it has recently been shown that sufficiently large Gaussian 
wave packets show a tendency to shrink in width as they evolve according to the 
time dependent Schr\"odinger-Newton equations. This leads 
to a {\it decrease of interference effects}, 
which possibly will be observable in near-future molecular 
interference experiments. It will be interesting to study the behaviour of such wave 
packets according to the present theory. 
We speculate that coherent superpositions of displaced wave packets 
({\it Schr\"odinger cat states}) will decay by giving rise to time dependent 
`potentials' ${\cal A}_t$ and ${\cal A}_\phi$, while attracting each other 
similar to corresponding classical matter distributions. -- 
If these remarks can be further substantiated, this should have some impact on 
further attempts to understand the ``collapse of the wave function'' or ``reduction of the 
wave packet'' in a consistent dynamical theory. 

\section{Concluding remarks}
A relativistic ${\cal U}(1)$ gauge theory has been presented which constitutes an 
intrinsically nonlinear extension of quantum mechanics or quantum field theory. 

Closest in spirit seems the work of Kibble and Randjbar-Daemi \cite{Kibble80} where 
such nonlinearities -- due to coupling the expectation of the quantum matter 
stress-energy tensor to classical general relativity or to making parameters of the 
theory state dependent -- have been 
discussed in a relativistic setting before. However, this has been reminiscent of a 
mean-field approximation. 
   
In distinction, based on the gauge principle, we have introduced two `potentials', 
${\cal A}_t$ and ${\cal A}_\phi$, which are {\it not independent} new fields but 
functionals that depend on the same field variables of the underlying (scalar or other) 
field theory as the wave functional $\Psi$. The 
relevant dynamical and constraint equations follow from a relativistic invariant 
action principle, postulated in Section\,3. Thus, if the `potentials' 
are eliminated, in principle, by solving the respective equations, a nonlinear theory in $\Psi$ 
necessarily results. 

We observe that in the absence of quantum matter, 
$\Psi =0$, the Eqs.\,(\ref{fieldeq}) and (\ref{Gauss}) that determine the `field strength' 
${\cal F}_{t\phi}$ -- and similarly in the (0+1)-dimensional limit -- have no time 
dependent solutions. Therefore, the `potentials' do not propagate independently of matter 
sources here.\footnote{This is due to the fact that the analogue of a magnetic 
field is missing for any underlying model based on a one-component field, see  
Eqs.\,(\ref{Lagrangean}) and (\ref{field}). The situation changes in the presence of internal 
symmetries, as discussed in the second of Refs.\,\cite{I06}.} 

We have shown that the essential homogeneity property holds, which is related to the 
representation of states by rays in Hilbert space. Thus, the Born rule can be applied, 
giving a probabilistic interpretation to $\Psi^*\Psi$ \cite{Kibble78,Iwo76,Weinberg89}. 
However, it   
breaks down, if the assumed underlying classical model contains dimensionfull parameters. 
In this case, a discussion in terms of the `charged' components of $\Psi$ is appropriate, 
which invites further interpretation.
  

Related to the presence of a fundamental lenght $l$ in the action, we have 
shown that in the zerodimensional limit the presented theory recovers the 
recently much studied Schr\"odinger-Newton equations, coupling Newtonian gravity 
to quantum mechanics \cite{Kibble80,Diosi84,Penrose98,Diosi05,Carlip06,Diosi06}. 
Thus, the proposed theory incorporates  
Newtonian gravity into quantum field theory: {\it unlike independent 
degrees of freedom coupled to matter in the usual way, gravity is encoded here into a 
universal nonlinear extension of quantum field theory}.   

In the future, the regularization of the theory and a  
perturbative scheme need to be worked out, in order to have control of its microscopic 
behaviour in situations where gravity is weak. Naturally, it will be interesting whether 
the presented ideas of a new {\it functional gauge symmetry} 
can be further generalized and what ensuing experimental predictions will be. 

\section*{Acknowledgement} 
It is a pleasure to thank H.-D.\,Doebner, G.A.\,Goldin, and C.\,Kiefer for discussions 
and correspondence during various stages of this work. The constructive suggestions 
from a referee have been highly appreciated.    

 

\end{document}